# MSM_pseudo_CS

### v1

@@ Abstract
Case-I interval-censored (current status) data from multistate systems are often encountered in cancer and other epidemiological studies. In this article, we focus on the problem of estimating state entry distribution and occupation probabilities, contingent on a preceding state occupation. This endeavor is particularly complex owing to the inherent challenge of the unavailability of directly observed counts of individuals at risk of transitioning from a state, due to the cross-sectional nature of the data. We propose two nonparametric approaches, one using the fractional at-risk set approach recently adopted in the right-censoring framework and the other a new estimator based on the ratio of marginal state occupation probabilities. Both estimation approaches utilize innovative applications of concepts from the competing risks paradigm. The finite-sample behavior of the proposed estimators is studied via extensive simulation studies where we show that the estimators based on severely censored current status data have good performance when compared with those based on complete data. We demonstrate the application of the two methods to analyze data from patients diagnosed with breast cancer.
@@ 1 Introduction
Multistate models are a popular type of survival model that allows for studying the progression of the disease using multivariate time-to-event data. In biomedical applications, situations arise in which patients progress reversibly or irreversibly through a succession of disease stages that describe their current condition. The patients usually begin from an initial disease state and move through one or more intermediate states before entering a final absorbing state. The disease stages correspond to the states of the multistate system. Moreover, in certain applications, the transition of patients through their multistate system may be subject to a severe form of (case-I) interval censoring often referred to as current status censoring (Jewell and van der Laan, 2003; Sun, 2006). Under such an observation scheme, the researcher is only restricted to information on the state occupied by the patient at a (single) random inspection time during the entire study period.
Current status data from multistate models often arise in epidemiological and biomedical research, as well as in observational studies where continuous monitoring of study participants is not feasible. With event history data from a multistate system, practitioners are often interested in estimating the probability of occupying a given state, at a specific time. Datta and Sundaram (2006) proposed marginal estimators for state occupation probabilities following the product integration approach of Aalen and Johansen (1978). Their nonparametric approach is valid without imposing the Markov assumption, and it has been further studied under different settings (Lan and Datta, 2010; Lan et al., 2017; Anyaso-Samuel et al., 2023).
In progressive multistate systems, where individuals navigate through successive states within a disease process, a pertinent question arises: “What is the probability of occupying a future state $k$, given a prior visit to state $j$?” We denote this conditional probability as $\Psi_{k|j}$. Furthermore, interest may extend to studying the probability of reaching a future state $k$ by time $t$, given a previous visit to state $j$, represented as $\Psi_{k|j}(t)$. Historically, these conditional probabilities have garnered less attention, as multistate systems are often assumed to adhere to Markovity. However, practitioners increasingly recognize the limitations of imposing a Markov assumption, which may be unverifiable and violated. Given a progressive multistate system, understanding the proportion of individuals on a unique path capable of occupying a future state becomes crucial. Our paper addresses this gap by focusing on the conditional estimation of future state occupation probabilities given prior occupation of a specific state, within the context of current status censoring.
In the context of right-censored data, Yang et al. (2023) recently proposed an estimator for the future state occupational probability conditional on a past state visit. Constructing such estimators is difficult since information on the transitions out of the non-absorbing states within the multistate system is only partially observed. With current status data, the problem becomes even more challenging since the path that individuals would have taken after they were observed is unknown. To tackle this problem of estimating $\Psi_{k|j}$ and $\Psi_{k|j}(t)$, we propose two estimation methods. The first is based on fractional at-risk sets. Here, we formulate the at-risk processes with weights defined by the fractional observation that a person contributes to the at-risk set of transitioning out of state $j$. We estimate this fractional weight as the probability that a patient will eventually reach state $j$ given that they were occupying a known state at a random inspection time. The idea for utilizing fractional weighted at-risk sets follows from the seminal work of Datta and Satten (2000) and the work of Yang et al. (2023). We begin the construction of the second estimator by providing insight that the conditional quantity $\Psi_{k|j}$ can be expressed as the ratio of two marginal state occupation probabilities for multistate models with a tree structure. This leads to the formulation of an estimator based on product limit estimators of state occupation probabilities (Datta and Sundaram, 2006). For both estimation methods, we do not impose the assumption that the multistate systems are Markov, but the methods require that the systems have a directed tree structure. Following our construction of estimators for the conditional future state occupation probabilities, we derive estimators for the distribution function of the state entry conditional on a prior occupation of a given state.
@@
The rest of the paper is organized as follows. In Section 2, we introduce the notation along with marginal nonparametric estimators for current status multistate data. In Section 3, we discuss the estimation of $\Psi_{k|j}$ based on the fractional weight approach. The alternate estimator based on the product-limit method is discussed in Section 4. In Section 5, we describe the estimation of pointwise confidence intervals for the estimators based on both approaches. Detailed simulation studies to evaluate the performance of the proposed estimators are presented in Section 6. In Section 7, we illustrate the application of both estimation methods in analyzing data from a breast cancer study. The paper concludes with a discussion in Section 8.
2 Nonparametric estimators for a multistate system
@@ 2.1 Notation and Convention
Consider a multistate system with a directed tracking structure, where $\mathcal{S}=\{0,1,\ldots,Q\}$ represents the finite state space. In this system, each individual $i=1,\ldots,n$ occupies a state denoted by $S_{i}(t)\in\mathcal{S}$ at time $t\geq 0$. Additionally, $C_{i}$ stands for the random inspection time for individual $i$. For an individual, the observed current-status data is represented as the pair $\{C_{i},S_{i}(C_{i})\}$, where $C_{i}$ is assumed to be independent of the multistate process $\{S_{i}(t),t\geq 0\}$, based on the random censoring assumption. To model the multistate current-status data, we adopt a framework where each state $j\in\mathcal{S}$ can be reached from the initial state 00 through a unique path: $0=s_{1}\rightarrow s_{2}\rightarrow\ldots\rightarrow s_{(j)+1}=j$, where $(j)$ denotes the number of steps it takes to reach $j$ from the root node 0.
Let $U_{jj^{\prime}}$ denote the (unobserved) transition time at which an individual transitions from $j$ to $j^{\prime}$. To estimate temporal functions of the multistate system, two essential functionals are the counting process denoted by $N_{jj^{\prime}}(t)$ and the at-risk process denoted by $Y_{j}(t)$. The counting process $N_{jj^{\prime}}(t)$ counts the number of transitions from state $j$ to $j^{\prime}$ in the time interval $[0,t)$. On the other hand, the at-risk process $Y_{j}(t)$ represents the number of individuals who are at risk of transitioning out of state $j$ at time $t$. These processes are expressed as follows
$Y_{j}(t)$ $=\sum^{n}_{i=1}\mathcal{I}(S_{i}(t-)=j),$
where $\mathcal{I}(\cdot)$ denotes the indicator function and $X(t-)$ is the state occupied just before time $t$. By the law of large numbers,
[…]
Using the theory of nonparametric regression, Datta and Sundaram (2006) proposed nonparametric estimators of $N_{jj^{\prime}}(t)$ and $Y_{j}(t)$ under this context of current status data. For $N_{jj^{\prime}}(t)$, they proposed a two-step estimation approach where first, isotonic regression via the pooled adjacent violators (PAV) algorithm (Barlow et al., 1972) is performed based on the pairs $(C_{i},\ \mathcal{I}(U_{i,jj^{\prime}}\leq C_{i}))$. Let $\widehat{N}^{P}_{jj^{\prime}}(t)$ denote the fitted responses from the isotonic regression. Then secondly, kernel smoothing (Mukerjee, 1988; Nadaraya, 1964; Watson, 1964) with a log-concave density, $K>0$ is performed to remove long flat parts of $\widehat{N}^{P}_{jj^{\prime}}(t)$. The resulting estimator of $N_{jj^{\prime}}(t)$ is given by
where $K_{h}(\cdot)=h^{-1}K(\cdot/h)$ with a bandwidth sequence $0<h=h(n)\downarrow 0$ chosen by the criteria specified by Wand and Jones (1995). Moreover, to estimate $\text{Pr}(X(t-)=j)$, which represents the probability limit of $n^{-1}Y_{j}(t)$, Datta and Sundaram (2006) employed the kernel smoothing approach. Accordingly, the estimation of $Y_{j}(t)$ is given by:
where it is assumed that $S_{i}(C_{i}-)=S_{i}(C_{i})$ with probability 1. The estimators outlined in (1) and (2) pave the way for feasible inference in multistate models with current status data. For example, Datta and Sundaram (2006) elaborates on the estimation of key marginal functionals, such as the transition probabilities and state occupation probabilities, employing the product limit approach described by Andersen et al. (1993).
This paper primarily proposes estimators for future state occupation probabilities and state entry distributions conditioned on a prior state visit. We employ the methodology described above to estimate the counting process and at-risk set, which serves as the foundational framework for constructing our proposed conditional estimators.
@@ 3.1 An illustration using a five-state model
To provide context for the aim of this study, let’s examine the progressive five-state illness-death model depicted in Figure 1. Our focus centers on estimating the probability of ever visiting state 3, given a prior stay in state 1 over an individual’s lifespan. We denote this conditional future occupation probability as $\Psi_{3|1}$. Estimating $\Psi_{3|1}$ poses a challenge due to incomplete observation of transition times. With current status data, we only have access to information about the state a patient occupies at a random inspection time.
0. Disease-free 1. Ill 2. Death without illness 3. Death after illness 4. Death from other causes State 0*
@@
Let $X_{ij}$ denote the (unobserved) indicator of the event that the $i$th individual would ever enter state $j$. In the typical current status framework, $X_{ij}$ will likely be unknown, making estimating the future entry states difficult. For instance, if the $i$th individual was observed to have $S_{i}(C_{i})=0$, then $X_{i1}$ will be unobserved. Suppose the $i$th individual transitioned to state 1, they will be eligible for the competing risks (state 3 and state 4), but we would not know this because $X_{i1}$ is not observed. To overcome the difficulty in estimating $\Psi_{3|1}$, first, we define a new state $0^{*}$ which is constructed by combining states 0, 1, and 2 (see Figure 1). Note that not everyone (i.e., those who transitioned to state 2) in state $0^{*}$ will be subject to the competing risks.
@@
[…]
where $\widehat{N}_{\ell\ell^{\prime}}(t)$ and $\widehat{Y}_{\ell}(t)$ are obtained by (1) and (2), respectively. The calculation of (3) does not require the Markov assumption since it pertains to the transition out of the root node. If $S_{i}(C_{i})=2$, then $\widehat{\phi}_{i1}(C_{i},S_{i}(C_{i}))=0$. For $S_{i}(C_{i})\in\{1,3,4\}$, indicating that the individual followed the $0\rightarrow 1$ path, $\widehat{\phi}_{i1}(C_{i},S_{i}(C_{i}))=1$.
@@
[…]
Here the superscript $[1]$ is used to distinguish this estimator from the alternate estimator discussed in Section 4. The estimators of the counting processes and the at-risk set in (4) are defined by
@@
Here, the estimation of the subdistribution function $\widehat{\Psi}^{[1]}_{3|1}(t)$ is achieved by constraining the integral’s bounds as defined in the expression provided in Equation 4. Therefore,
[…]
@@ 3.2 Extension to general multistate model with a tree structure
Next, we expand the explanation of the estimator for $\Psi_{k|j}=P(X_{k}=1~{}|~{}X_{j}=1)$ to encompass a broader multistate model context. The estimator (based on fractional weights) of the probability of occupying state $k$ given a prior visit to state $j$ is denoted as $\widehat{\Psi}^{[1]}_{k|j}$. The estimation of ${\Psi}_{k|j}$ relies on the assumption that state $j$ lies on the unique path connecting states 00 and $k$; otherwise, ${\Psi}_{k|j}=0$. Let $m$ represent the number of transitions needed to move from state 00 to state $k$. Constructing the estimator for $\Psi_{k|j}$ follows a recursive approach based on $m$.
@@
where $\tilde{k}$ is the state just before state $k$ on the path from 00 to $k$. From this recursive formula, $\widehat{\Psi}^{[1]}_{\tilde{k}|j}$ is available since the distance from state 0 to state $\tilde{k}$ is less than $m$. To construct the estimator $\widehat{\Psi}^{[1]}_{k|\tilde{k}}$, we utilize the idea of estimating transition probabilities in the competing risk framework and the concept of “pooling of states”. Thus, we have
@@
Further, we describe the estimation of the entry distribution for state $k$ given a previous occupancy of state $j$, denoted as $\widehat{F}_{k|j}(t)$. Let $\mathcal{S}^{k}$ denote the set of states $\ell$ such that state $k$ is on the path from state 0 to state $\ell$. Then, $\widehat{F}_{k|j}(t)$ can be expressed as:
$\widehat{F}_{k|j}(t)=\frac{\sum_{\ell\in\mathcal{S}^{k}}\widehat{\Psi}^{[1]}_{\ell|j}(t)}{\sum_{\ell\in\mathcal{S}^{k}}\widehat{\Psi}^{[1]}_{\ell|j}},$ (8)
where $\widehat{\Psi}^{[1]}_{k|j}(t)=\widehat{\Psi}^{[1]}_{\widetilde{k}|j}(t)\cdot\widehat{\Psi}^{[1]}_{k|\widetilde{k}}(t)$, as defined in Eqn. (6). Note that, $\widehat{F}_{k|j}(t)$ is the normalized sum of estimated future occupation probabilities (conditioned on a prior visit to state $j$) of state $k$ and all other states that come after state $k$ in the progressive system.
For a practical illustration, consider a seven-state multistate system (refer to Figure 4). Suppose we aim to estimate the probability of occupying state 5 given a prior visit to state 1, denoted by $\Psi_{5|1}$. Utilizing Eqn. (6), in this case, we have $k=5$, $j=1$, and $\widetilde{k}=3$. Consequently, our goal is to compute $\widehat{\Psi}^{[1]}_{5|1}=\widehat{\Psi}^{[1]}_{3|1}\cdot\widehat{\Psi}^{[1]}_{5|3}$. Following the recursive calculation, $\widehat{\Psi}^{[1]}_{3|1}$ is already available. To calculate $\widehat{\Psi}^{[1]}_{5|3}$ using (7), define the artificial state $0^{*}=\{0,1,2,3,4\}$. Moreover, for the non-trivial case, the fractional weight $\widehat{\phi}_{i3}$ is obtained by estimating the probability of transitioning from another artificial state defined by pooling states $\{0,1,2\}$, to state 3.

### v2

# Nonparametric estimation of a state entry time distribution conditional on a “past” state occupation in a progressive multistate model with current status data

@@ Abstract
Case-I interval-censored (current status) data from multistate systems are often encountered in biomedical and epidemiological studies. In this article, we focus on the problem of estimating state entry distribution and occupation probabilities, contingent on a preceding state occupation. This endeavor is particularly complex owing to the inherent challenge of the unavailability of directly observed counts of individuals at risk of transitioning from a state, due to severe interval censoring. We propose two nonparametric approaches, one using the fractional at-risk set approach recently adopted in the right-censoring framework and the other a new estimator based on the ratio of marginal state occupation probabilities. Both estimation approaches utilize innovative applications of concepts from the competing risks paradigm. The finite-sample behavior of the proposed estimators is studied via extensive simulation studies where we show that the estimators based on severely censored current status data have good performance when compared with those based on complete data. We demonstrate the application of the two methods to analyze data from patients diagnosed with breast cancer.
@@ 1 Introduction
Multistate models are a foundational framework for analyzing disease progression with multivariate time-to-event data in biomedical research. Individuals typically start in an initial state and move through intermediate disease states before possibly entering an absorbing state. In many settings, however, continuous follow-up is infeasible and the process is observed only at a single random inspection time per individual (“current status” or case-I interval censoring), so that for each individual one records only the state occupied at that inspection [29, 14, 28].
In progressive systems with a directed tree structure, a clinically pertinent question is: what is the probability of ever occupying a given state $k$, conditional on a prior visit to state $j$? We denote this marginal (population-level) probability by $\Psi_{k\mid j}$, and its time-specific analogue by $\Psi_{k\mid j}(t)$. These quantities summarize the proportion of individuals who, among those who ever visit $j$, ultimately (or by time $t$) occupy $k$ on the unique path from the root to $k$. Estimation of such marginal-conditional probabilities can inform prognosis and planning after intermediate milestones. For example, risk of distant metastasis after loco-regional recurrence in oncology or risk of a second remission after a first remission in hematologic malignancies [25, 17, 30, 22].
For right-censored event-history data from progressive systems, estimators of state-occupation probabilities conditional on prior state occupation have recently appeared [34]. Under current-status observation, the problem is tougher: one never observes the future path after the single inspection, and even transitions out of nonabsorbing states are only partially observed. This paper develops two nonparametric estimators of $\Psi_{k\mid j}$ and $\Psi_{k\mid j}(t)$ for progressive multistate models observed under current status, without imposing the Markov property. The first estimator builds fractional at-risk sets by weighting each individual’s contribution to the at-risk set of transitioning out of state $j$ by the probability that the individual would ever reach $j$, adapting ideas of fractional weighting from Datta and Satten[8] and connections to Yang et al. [34] The second estimator exploits the identity of $\Psi_{k\mid j}$ as a ratio of marginal occupation probabilities in tree-structured systems, leading to a product-limit–type construction based on multistate state-occupation estimators [10]. We also derive estimators of the distribution of the entry time into $k$ given prior occupation of $j$.
@@
The rest of the paper is organized as follows. Section 2 formalizes the single-inspection multistate design and gives biomedical examples where such data arise. Sections 3 and 4 develop the fractional-risk and product-limit estimators, respectively. In Section 5, we describe the estimation of pointwise confidence intervals for the estimators based on both approaches. Section 6 discusses testing of covariate effects via pseudo-value regression. Detailed simulation studies to evaluate the performance of the proposed estimators are presented in Section 7. In Section 8, we illustrate the application of both estimation methods in analyzing data from a breast cancer study. The paper concludes with a discussion in Section 9.
2 Study Design: Single-Inspection Multistate Sampling
2.1 What is observed under multistate current status?
Consider a progressive multistate process $\{S_{i}(t):t\geq 0\}$ on a directed tree $\mathcal{S}=\{0,1,\ldots,Q\}$ with unique root-to-node paths. Under current-status observation, subject $i$ is inspected once at a random time $C_{i}$ and only the pair $\{C_{i},S_{i}(C_{i})\}$ is recorded. The inspection times $\{C_{i}\}$ are independent of the underlying multistate process, based on the random censoring assumption. No transition times or future trajectories after $C_{i}$ are observed. Thus, the data consist of a single cross-sectional snapshot of the state occupancy across heterogeneous times since a common origin (e.g., time since diagnosis, surgery, transplantation, or age).
This observation scheme is stricter than right-censoring but is practically motivated whenever repeated follow-up is infeasible, expensive, or ethically constrained (e.g., one-time biospecimen collection).
2.2 How common is the design in practice?
We acknowledge that multistate current-status data are not as ubiquitous as right-censored multistate cohorts in public repositories. Nevertheless, single-inspection designs arise in several biomedical domains, and importantly, they produce multistate current-status information:
  1. (a)
    Two-state (classical) current status in screening and serosurveys. Cross-sectional surveys that ascertain whether an event has occurred by age or time-since-exposure are textbook examples of current-status designs in epidemiology and screening [28].
  2. (b)
    Progressive infection staging from a single biospecimen. HIV recency assays classify cross-sectional specimens into “recent” versus “long-standing” infection to support incidence surveillance; individuals can be viewed as moving from susceptible $\to$ recent infection $\to$ established infection along a progressive path observed at one time point [15, 11].
  3. (c)
    Serology-defined multistate profiles with branching. Hepatitis B screening panels (HBsAg, anti-HBc, anti-HBs) classify single-visit laboratory results into mutually exclusive clinical states: susceptible, immune due to vaccination, immune from past infection, chronic infection; forming a progressive tree with extended branches observable at one inspection (e.g., by age) [26].
  4. (d)
    Staged chronic disease from a single clinical exam. Periodontitis is routinely staged in cross-sectional surveillance, producing ordinal disease states (Stage I–IV) on a progressive path observed at one dental exam; recent methodological work is explicitly motivated by cross-sectional multistate current-status periodontal data [4]. Diabetic retinopathy screening programs similarly classify a single fundus image into graded states used for referral and surveillance, again yielding a progressive multistate snapshot [24].
Although multistate current-status datasets with extended branching are less common in public repositories than right-censored cohorts, the design often arises in practice as indicated above. Because publicly available examples with rich branching are scarce, for illustration, researchers often emulate a single-inspection design from an existing right-censored dataset by masking follow-up to one random inspection per subject, consistent with the random inspection-time assumption.
@@ 2.3 Notation and Convention
For each state $k\in\mathcal{S}$, let $\mathcal{P}(k)=\{0=s_{1}\rightarrow s_{2}\rightarrow\cdots\rightarrow s_{m}=k\}$ denote the unique directed path from the root node 0 to state $k$, where $m$ is the number of steps required to reach $k$. A state $j$ is said to lie on the path to $k$ if $j\in\mathcal{P}(k)$ and $j\neq k$.
Let $U_{i,jj^{\prime}}$ denote the (unobserved) transition time at which individual $i$ transitions from $j$ to $j^{\prime}$. Also, let $X_{ij}$ be the (unobserved) indicator that individual $i$ would eventually visit state $j$. The main estimands of interest, $\Psi_{k\mid j}$ and its time-specific version $\Psi_{k\mid j}(t)$, are defined for states $j\in\mathcal{P}(k)$ and represent, respectively, the marginal probability of ever occupying state $k$ given a prior visit to $j$, and the probability of occupying state $k$ by time $t$ given prior occupation of $j$. These quantities are defined at the population level and are identifiable from single-inspection data under the assumptions described.
@@ 3.1 An illustration using a five-state model
To provide context for the aim of this study, let’s examine the progressive five-state illness-death model depicted in Figure 1. Here, state occupation would be determined through medical examinations. Our focus centers on estimating the probability of occupying state 3, given a prior visit to state 1 over an individual’s lifespan. We denote this conditional occupation probability as $\Psi_{3|1}$. Estimating $\Psi_{3|1}$ poses a challenge due to incomplete observation of transition times. With current status data, we only have access to information about the state a patient occupies at a random inspection time.
0. Disease-free 1. Ill 2. Death without illness 3. Death due to illness 4. Death from other causes State 0*
@@
In the typical current status framework, $X_{ij}$ will likely be unknown, making inference on a later state difficult. For instance, if the $i$th individual was observed to have $S_{i}(C_{i})=0$, then $X_{i1}$ will be unobserved. Suppose the $i$th individual transitioned to state 1, they will be eligible for the competing risks (state 3 and state 4), but we would not know this because $X_{i1}$ is not observed. To overcome the difficulty in estimating $\Psi_{3|1}$, first, we define a new state $0^{*}$ which is constructed by combining states 0, 1, and 2 (see Figure 1). Note that not everyone (i.e., those who transitioned to state 2) in state $0^{*}$ will be subject to the competing risks.
@@
$\widehat{\phi}_{i1}(C_{i},S_{i}(C_{i}))=\int_{C_{i}}^{\infty}\left\{\prod_{(C_{i},u)}\left[1-\frac{d\widehat{N}_{01}(v)+d\widehat{N}_{02}(v)}{\widehat{Y}_{0}(v)}\right]\right\}\frac{d\widehat{N}_{01}(u)}{\widehat{Y}_{0}(u)}.$ (1)
Here, $\widehat{N}_{\ell\ell^{\prime}}(t)$ is the marginal estimator of the counting process $N_{\ell\ell^{\prime}}(t)$, which counts the transitions from state $\ell$ to state $\ell^{\prime}$ in the time interval $[0,t)$, and $\widehat{Y}_{\ell}(t)$ is the estimator of the at-risk process $Y_{\ell}(t)$, representing the number of individuals at risk of transitioning out of state $\ell$ at time $t$. The details of the construction of these estimators via nonparametric regression are provided in Appendix A (ref: (10)-(11)).
Importantly, the calculation of (1) does not rely on the Markov assumption, as it only concerns the transition from the root node (state 0). If $S_{i}(C_{i})=2$, then $\widehat{\phi}_{i1}(C_{i},S_{i}(C_{i}))=0$. For $S_{i}(C_{i})\in\{1,3,4\}$, indicating that the individual followed the $0\rightarrow 1$ path, $\widehat{\phi}_{i1}(C_{i},S_{i}(C_{i}))=1$.
@@
$\widehat{\Psi}^{[1]}_{3|1}=\int^{\infty}_{0}\left\{\prod_{(0,u)}\left[1-\frac{d\widehat{N}_{3|1}(v)+d\widehat{N}_{4|1}(v)}{\widehat{Y}_{0^{*}|1}(v)}\right]\right\}\frac{d\widehat{N}_{3|1}(u)}{\widehat{Y}_{0^{*}|1}(u)}.$ (2)
Here the superscript $[1]$ is used to distinguish this estimator from the alternate estimator discussed in Section 4. To conduct the nonparametric regression, let $K>0$ represent a log-concave (differentiable) density, and let $0<h=h(n)\downarrow 0$ be a bandwidth sequence. The estimators of the counting processes and the at-risk set in (2) are defined by
@@
Here, the estimation of the subdistribution function ${\Psi}^{[1]}_{3|1}(t)$ is achieved by constraining the integral’s bounds as defined in the expression provided in (2). Therefore,
$\widehat{\Psi}^{[1]}_{3|1}(t)=\int^{t}_{0}\left\{\prod_{(0,u)}\left[1-\frac{d\widehat{N}_{3|1}(v)+d\widehat{N}_{4|1}(v)}{\widehat{Y}_{0^{*}|1}(v)}\right]\right\}\frac{d\widehat{N}_{3|1}(u)}{\widehat{Y}_{0^{*}|1}(u)}.$ (3)
@@ 3.2 Extension to general multistate model with a tree structure
Next, we expand the explanation of the estimator for $\Psi_{k|j}=P(X_{k}=1~|~X_{j}=1)$ to encompass a broader multistate model context. The estimator (based on fractional weights) of the probability of occupying state $k$ given a prior visit to state $j$ is denoted as $\widehat{\Psi}^{[1]}_{k|j}$. The estimation of ${\Psi}_{k|j}$ relies on the assumption that state $j$ lies on the unique path connecting states 0 and $k$; otherwise, ${\Psi}_{k|j}=0$. Let $m$ represent the number of transitions needed to move from state 0 to state $k$. Constructing the estimator for $\Psi_{k|j}$ follows a recursive approach based on $m$.
@@
where $\tilde{k}$ is the state just before state $k$ on the path from 0 to $k$. From this recursive formula, $\widehat{\Psi}^{[1]}_{\tilde{k}|j}$ is available since the distance from state 0 to state $\tilde{k}$ is less than $m$. To construct the estimator $\widehat{\Psi}^{[1]}_{k|\tilde{k}}$, we utilize the idea of estimating transition probabilities in the competing risk framework and the concept of “pooling of states”. Thus, we have
@@
Let $T_{k}$ denote the random time at which state $k$ is reached. Estimation of the subdistribution function, $\Psi_{k|j}(t)=\text{Pr}(T_{k}\leq t,X_{k}=1\mid X_{j}=1)$ also follows (4). That is; $\widehat{\Psi}^{[1]}_{k|j}(t)=\widehat{\Psi}^{[1]}_{\tilde{k}|j}(t)\cdot\widehat{\Psi}^{[1]}_{k|\tilde{k}}(t)$.
We are also interested in the entry distribution for state $k$ given a previous occupancy of state $j$, expressed as
${F}_{k|j}(t)=\text{Pr}(T_{k}\leq t\mid X_{k}=1,\ X_{j}=1),$
that is, the cumulative distribution function of $T_{k}$ among individuals who eventually reach $k$ after having reached $j$. By conditioning on $X_{j}=1$, we have
$\begin{aligned} {F}_{k|j}(t)&=\frac{\text{Pr}(T_{k}\leq t,\ X_{k}=1\mid X_{j}=1)}{\text{Pr}(\ X_{k}=1\mid X_{j}=1)}=\frac{\Psi_{k|j}(t)}{\Psi_{k|j}}\end{aligned}$
Then, the estimator is obtained by
${F}_{k|j}(t)=\frac{\widehat{\Psi}_{k|j}(t)}{\widehat{\Psi}_{k|j}}.$ (6)
For a practical illustration, consider a seven-state multistate system (refer to Figure 2). Suppose we aim to estimate the probability of occupying state 5 given a prior visit to state 1, denoted by $\Psi_{5|1}$. Utilizing (4), in this case, we have $k=5$, $j=1$, and $\widetilde{k}=3$. Consequently, our goal is to compute $\widehat{\Psi}^{[1]}_{5|1}=\widehat{\Psi}^{[1]}_{3|1}\cdot\widehat{\Psi}^{[1]}_{5|3}$. Following the recursive calculation, $\widehat{\Psi}^{[1]}_{3|1}$ is already available. To calculate $\widehat{\Psi}^{[1]}_{5|3}$ using (5), define the artificial state $0^{*}=\{0,1,2,3,4\}$. Moreover, for the non-trivial case, the fractional weight $\widehat{\phi}_{i3}$ is obtained by estimating the probability of transitioning from another artificial state defined by pooling states $\{0,1,2\}$, to state 3.